OpenStudy (anonymous):

Sorry to bother but can someone help me understand this thank you so much.How many kilograms (kg) of pure salt must be added to 8 kg of a 30% solution to obtain a 35% solution? Round to the nearest tenth.

OpenStudy (anonymous):

at the moment in solution 30% is pure salt 0.3*8 = 2.4kg of pure salt initially lets say we add x kg of salt now we have 2.4 + x kg of pure salt in 8+x kg of total solution we want the ratio pure:solution to be 35:100 we get the following equation because \[\frac{pure}{total} = \frac{35}{100}\] \[\frac{2.4+ x}{8+x} = \frac{35}{100}\] \[2.4 + x = 0.35(8+x)\] \[2.4 + x = 2.8 + 0.35x\] \[x - 0.35x = 2.8 - 2.4\] \[0.65x = 0.4\] \[x = \frac{0.4}{0.65} = 0.615...\] so 0.6 kg to the nearest tenth of a kilogram

OpenStudy (anonymous):

I understand it much better now thank you so much I was getting confused thank you

OpenStudy (anonymous):

:) happy to help